OpenStudy (anonymous):

Find all numbers x for which \[\frac{1}{x}+\frac{1}{1-x} > 0\]

OpenStudy (anonymous):

hard? :)

myininaya (myininaya):

\[\frac{1-x+x}{x(1-x)} >0\] \[\frac{1}{x(1-x)} >0\] \[f(x)=\frac{1}{x(1-x)} \text{ is undefined at x=0 & x=1 }\] -----|----------|----- 0 1 We need to choose a number from the following intervals: (-inf,0) ; (0,1) ; (1,inf) /\ /\ /\ | | | -1 .5 2 f(-1)=- f(.5)=+ f(2)=- We want to know where f>0 so f>0 on the interval (0,1)

myininaya (myininaya):

Nope I just wanted to explain it :)

OpenStudy (anonymous):

thanks I was solving this one but I couldn't figure out why my answer was different from the given answer

myininaya (myininaya):

did you try to clear the denominators?

OpenStudy (anonymous):

I just thought that it was postive if both numerator and denominator are positive, or both are negative

myininaya (myininaya):

The bad thing about clearing the denominators is that x is not always greater than 0 and 1-x is not always positive so we can't just simply multiply both sides by x and (1-x)

OpenStudy (anonymous):

right.

myininaya (myininaya):

Did you right it as one fraction like i did?

myininaya (myininaya):

Or did you look at the fractions separately?

myininaya (myininaya):

right=write my bad